OpenStudy (anonymous):

can you help me find the the equlvalent expression for 6/13 with a denominator of 78a 6/13^(6a/6a) 36a/78a

jimthompson5910 (jim_thompson5910):

Notice how 78/13 = 6 So we need to multiply both the top and bottom of the fraction by 6a to get the denominator equal to 78a The numerator then becomes 6*6a = 36a So the final fraction is \[\large \frac{36a}{78a}\]